OpenStudy (anonymous):

Pat is 24 years older than his son Patrick. In 2 years Pat will be twice as old as Patrick. How old are they now?

OpenStudy (s):

Patrick = x Pat = x+24 2*(x+2) = x+24+2 2x+4 = x+26 2x-x = 26-4 x = 22 So now, Patrick is 22 and Pat is 22+24=46 In 2 years, Patrick will be 22+2=24 and Pat will be 46+2=48.. which is twice as much as 24.